# Symmetry in Neural Network Parameter Spaces

### v2

@@ 1 Introduction
Despite their remarkable empirical success, neural networks remain not fully understood from a theoretical standpoint. Designing effective architectures often relies on heuristics, and training outcomes can be highly sensitive to initialization and optimization details. The rapid growth of model sizes—exemplified by the recent explosion in the size of language and other generative models (Kaplan et al., 2020)—further complicates efforts to characterize training dynamics and generalization behavior. While theoretical work has clarified aspects of overparameterization (Belkin et al., 2019; Neyshabur et al., 2018) and implicit regularization in gradient descent (Soudry et al., 2018), we still lack a unified understanding of how the structure of parameter space shapes the loss landscape and affects optimization.
@@
This survey proposes parameter space symmetry as a valuable, underexplored perspective for understanding neural networks. One of the most important consequences of parameter space symmetry is that it induces nontrivial structure in the level sets of the loss function. In overparameterized networks, where many parameter configurations can yield the same function output, the loss landscape often contains minima with high dimension (Cooper, 2018). Symmetry accounts for much of this degeneracy by mapping parameters within a level set without changing the network’s function. For example, permuting neurons in a hidden layer leaves the function unchanged but yields a different point in parameter space (Figure 1). Understanding these equivalence classes is essential for analyzing generalization, optimization dynamics, and the loss landscape in deep learning.
While symmetry in data spaces has been central to geometric deep learning and equivariant models (Bronstein et al., 2021), symmetry in parameter space has only recently begun to receive sustained attention. Recent work explores parameter symmetry in diverse contexts, including loss landscapes and mode connectivity (Garipov et al., 2018; Draxler et al., 2018), conserved quantities and training dynamics (Kunin et al., 2021; Tanaka and Kunin, 2021), symmetry-based optimization and model averaging (Entezari et al., 2022; Ainsworth et al., 2023), and symmetry-aware sampling in Bayesian inference (Wiese et al., 2023). This survey aims to define symmetries in neural network parameter spaces, unify existing theoretical and algorithmic perspectives, and highlight their relevance to learning theory, optimization, and model analysis. We hope that a clearer understanding of parameter space symmetries will lead to more principled approaches in deep learning theory and practice.
@@ 2.1 Loss-Invariant Parameter Transformations and Symmetry
Let $\Theta$ be the space of parameters and $\mathcal{D}$ be the space of data. In supervised learning, $\mathcal{D}$ is decomposed into $\mathcal{D}_{\text{input}}\times\mathcal{D}_{\text{target}}$, where $\mathcal{D}_{\text{input}}$ and $\mathcal{D}_{\text{target}}$ correspond to the space of inputs and target labels, respectively. Let $f\colon\Theta\times\mathcal{D}_{\text{input}}\to\mathcal{D}_{\text{target}}$ be a neural network function, and $c\colon\mathcal{D}_{\text{target}}\times\mathcal{D}_{\text{target}}\to{\mathbb{R}}$ be a function that measures the discrepancy between the neural network’s prediction and the ground truth label. The loss function $L\colon\Theta\times\mathcal{D}\to{\mathbb{R}}$ is the composition of $f$ and $c$. With occasional exceptions (Bassey et al., 2021; Huang et al., 2018), the parameter space of a neural network is typically a real vector space.
@@
The set of all such transformations forms a group under function composition. To see this, we first examine its structure and properties. Viewing the transformations as functions, we are able to define the composition of two transformations. The composition of two loss-invariant transformations results in another loss-invariant transformation. Since function compositions are associative, the composition of these transformations are associative. Additionally, the identity transformation is always loss preserving, therefore included in our set. Finally, since each transformation is bijective, an inverse exists for every element in the set. The properties of the set of symmetry transformations under composition satisfy the definition of a group.
@@
Following this definition, the set of all loss-invariant and bijective transformations, which we denote by $G_{\Theta,L}$, forms a group. In practice, we typically consider specific subgroups of $G_{\Theta,L}$, since the full group is often difficult to characterize.
@@ 2.2 Functional Neural Network Symmetry
Parameter space symmetry can be defined in various ways, depending on the transformation preserved, i.e. the neural function or the loss function, and the scope of the data considered, which can range from all possible data, subsets of data, or a particular distribution (Figure 5).
@@ 2.2.1 Examples: Symmetries in Common Neural Network Components
Often appearing in equivariant neural networks (Weiler et al., 2018; Weiler and Cesa, 2019), a radial rescaling activation $\sigma\colon{\mathbb{R}}^{h}\to{\mathbb{R}}^{h}$ has the form $\sigma(z)=f(\|z\|)z$ for some function $f\colon{\mathbb{R}}\to{\mathbb{R}}$. A two-layer neural network $f(W_{2},W_{1},X)=W_{2}\sigma(W_{1}X)$ with a radial rescaling activation $\sigma$ has an $\text{O}_{h}({\mathbb{R}})$ symmetry, acting on $\Theta$ by $g\cdot(W_{2},W_{1})=(W_{2}g^{-1},gW_{1})$, for $g\in\text{O}_{h}({\mathbb{R}})$.
@@
Example 2.6 (Batch Normalization (Ioffe and Szegedy, 2015)).
Batchnorm standardizes inputs of a layer across a mini-batch: $\text{BN}(Z)=\frac{Z-\text{E}[Z]}{\sqrt{\text{Var}[Z]}}$, where the expectation and variance are computed row-wise over $Z$. Assuming $Z$ is a linear feature $Z=WX$ where $W\in\Theta={\mathbb{R}}^{m\times n}$ and $X\in{\mathbb{R}}^{n\times k}$, then the batchnorm function has a positive scaling symmetry. Batchnorm is equivariant with respect to the group ${\mathbb{R}}_{>0}^{m}$, which consists of diagonal matrix with positive diagonal entries, acting on $\Theta$ by $g\cdot W=gW$, for $g\in{\mathbb{R}}_{>0}^{m}$ (Kunin et al., 2021). Other normalization layers, such as layer normalization (Ba et al., 2016), group normalization (Wu and He, 2018), and weight normalization (Salimans and Kingma, 2016), often have scaling symmetry of similar forms.
@@
Many architectures have parameter symmetries given by a symmetric group. This symmetry can often be interpreted as permutation of neurons or components along with their associated weights, as shown in Examples 2.7, 2.8 and Figure 3(c).
@@
Example 2.9 (Radial basis function networks (Broomhead and Lowe, 1988)).
@@ 2.2.2 Example: Symmetries in Transformers
2.3 More General Definitions of Symmetry
@@ 2.4 Parameter Identifiability and Completeness of Symmetry
Having introduced a range of examples highlighting that symmetry is ubiquitous in neural network parameter spaces, we now turn to the fundamental question of whether known parameter symmetries account for all the ways in which parameters can represent the same function. This question concerns parameter identifiability—the extent to which a neural network’s function determines its parameters, which is important in interpreting and comparing trained networks (Ran and Hu, 2017). This section formalizes the relationship between symmetry and identifiability using the realization map from parameters to functions, and surveys results on when known symmetries are sufficient to characterize all such redundancy.
Formally, we consider the realization map $\rho\colon\Theta\to\mathcal{F},\text{where }\mathcal{F}=\{f\colon\mathcal{D}_{\text{input}}\to\mathcal{D}_{\text{output}}\}$, which associates each parameter $\theta\in\Theta$ with the function $\rho(\theta)\in\mathcal{F}$ it defines (Grigsby et al., 2022). The fiber of a function $f$ under $\rho$, denoted $\rho^{-1}(f):=\{\theta\in\Theta~|~\rho(\theta)=f\}$, is the set of all parameters that map to the same function $f$. In many architectures, $\rho$ is not injective, and fibers contain multiple distinct parameters that define the same function. This implies that parameters are, in general, not uniquely determined by the function alone.
@@
For several architectures, such completeness has been established. For example, Gaussian radial basis function networks are known to be identifiable up to permutation symmetry, indicating that permutation symmetry is complete for these architectures (Kůrková and Neruda, 1994). For feedforward networks with tanh activations, if two networks represent the same function, then they have the same architecture and parameters up to permutation and sign-flip symmetries, under a known set of conditions (Sussmann, 1992; Fefferman and Markel, 1993). Similar completeness results have also been established for complex-valued tanh networks (Nitta, 2003; Kobayashi, 2010), tanh recurrent neural networks (Kimura, 2002), networks with activation functions $\sigma$ satisfying $\sigma(0)=0,\sigma^{\prime}(0)\neq 0$, and $\sigma^{\prime\prime}(0)=0$ (Albertini and Sontag, 1993), and networks with asymptotically constant activation functions (Kůrková and Kainen, 1994).
@@
The identifiability of ReLU networks is less straightforward. In particular, ReLU networks can contain hidden symmetries—transformations that preserve the realized function but not captured by permutation and positive scaling symmetries. Grigsby et al. (2023) describe several mechanisms through which hidden symmetries can arise. For example, for a single ReLU neuron with scalar input $x$ and realization map $\rho\colon(a,b)\mapsto(x\mapsto\text{ReLU}(ax+b))$, the fiber of the constant function 0, $\rho^{-1}(0)$, contains $(0,0)$ and (0,1)(0,-1), which are not related by rescaling and permutation (Grigsby et al., 2022). To formalize this variability, Grigsby et al. (2022) introduce the functional dimension, defined as the rank of the Jacobian of the realization map $\rho$, which reflects the local dimensionality of the set of functions realized near a given parameter. They show that the functional dimension may not be constant even within the same fiber. This suggest a complicated structure of symmetries in ReLU networks.
@@
Other recent work investigates when ReLU networks are identifiable up to rescaling and permutation symmetries. (Petzka et al., 2020) prove that excluding a degenerate case where two neurons have identical zero hyperplanes, two-layer ReLU networks do not have additional symmetries besides permutation and positive scaling. (Rolnick and Kording, 2020) prove that under the assumption that networks within each linear region compute the same linear function, boundaries between linear regions uniquely determine parameters up to permutation and scaling. For ReLU networks with either non-increasing widths (Bui Thi Mai and Lampert, 2020) or all layers wider than input (Grigsby et al., 2023), a positive measure subset of the parameters have no symmetries besides permutation and positive scaling (hidden symmetries). Grigsby et al. (2023) also empirically show that the probability of having no hidden symmetry increases with input and layer width and decreases with depth. (Bona-Pellissier et al., 2023) provide sufficient conditions under which two ReLU network are identical up to permutation and positive scaling if their output agrees on a subset of the input space.
@@
The identifiability of neural networks that are polynomial in their parameters can often be analyzed using tools from algebraic geometry (Marchetti et al., 2025). A notable example is multi-layer perceptrons with polynomial activation functions, or polynomial neural networks. Kileel et al. (2019) conjecture that permutation and scaling are the only symmetries in generic polynomial neural networks, which has been proven in the monomial case (Finkel et al., 2024). For more general polynomial activation functions, Shahverdi et al. (2025) show that there are only finitely many data-independent parameter symmetries. Moreover, they prove that convolutional neural networks with polynomial activations generically have no nontrivial data-independent parameter symmetries. Another example of a polynomial function is lightening self-attention $(W^{Q},W^{K},W^{V})\mapsto QW^{Q}(KW^{K})^{T}VW^{V}$, where the softmax normalization is omitted from the standard attention mechanism (Example 2.10). (Henry et al., 2025) prove that in a deep network composed of lightning self-attention layers, the only symmetries are (1) scaling $W_{Q}W_{K}^{T}$ and $W_{V}$ by a constant, (2) scaling $W_{Q}$ and $W_{K}$ by an invertible matrix, and (3) scaling $W_{V}$ of one layer and $W_{Q},W_{K}$ in the next layer by an invertible matrix. They conjecture and verify numerically that adding back the softmax breaks only the first of these symmetries.
@@ 3.3 Removing Symmetry: Model Compression and Reduced Search Space
While symmetry enriches the loss landscape by introducing structure and multiplicity, removing symmetry reveals a simpler geometry—one that is more compact to represent, and often more amenable to analysis. Symmetry removal effectively collapses each symmetry orbit to a single representative. This operation alters the shape of the loss landscape—not by changing its values, but by restricting attention to a lower-dimensional quotient space where each point represents a unique function. As a result, the loss landscape becomes less redundant, more compact, and easier to sample from and optimize.
One major motivation for removing symmetry is to reduce the dimensionality of the parameter space without altering the function space. This operation, often known as model compression, is of critical importance in lowering storage cost, improving inference speed, and efficient deployment of machine learning models. Although most production pipelines use pruning, quantization, or distillation, symmetry-aware reductions offer a function-preserving complement that can be combined with these methods. Every symmetry implies a set of directions in which the loss is constant, stretching level sets into high-dimensional manifolds. Quotienting out these directions yields a more compact version of the loss landscape, where each point corresponds to a unique function. In the context of model compression, this means we can represent networks more efficiently. For example, Ganev et al. (2022) compress radial neural networks by factoring out orthogonal symmetries, and Sourek et al. (2021) exploit structural symmetries in computational graphs to merge redundant nodes. These methods reduce storage costs and inference time, without affecting expressivity or accuracy.
@@
In Bayesian neural networks, removing symmetry reduces the effective search space by collapsing functionally equivalent modes in the posterior (Lim et al., 2024b). In high-dimensional models like neural networks, symmetries can cause posterior distributions to become multi-modal in a functionally redundant way. This impairs MCMC mixing and complicates interpretation. By mapping parameters to a canonical representative—effectively quotienting out permutation or sign-flip symmetries—Wiese et al. (2023) and Laurent et al. (2024) show that it is possible to explore a more informative and diverse set of functionally distinct solutions. (Xiao et al., 2023) shows that removing permutation symmetry leads to a compact representation that helps directly compare trained BNNs across sampling methods.
Finally, removing symmetry alters the curvature and critical point structure of the landscape, with direct implications for optimization. Symmetry-induced flat directions often give rise to plateaus or saddle points that slow down training. By projecting the loss landscape onto a symmetry-reduced space, these degenerate directions are removed, yielding a more strongly convex surface. Leake and Vishnoi (2021) show that continuous symmetries can be used to construct convex polytopes, allowing nonconvex optimization problems to be reformulated as linear optimization problems over polytopes, which are computationally more tractable.

### v3

@@ 1 Introduction
Despite their remarkable empirical success, neural networks remain not fully understood from a theoretical standpoint. Designing effective architectures often relies on heuristics, and training outcomes can be highly sensitive to initialization and optimization details. The rapid growth of model sizes—exemplified by the recent explosion in the size of language and other generative models (Kaplan et al., 2020)—further complicates efforts to characterize training dynamics and generalization behavior. While theoretical work has clarified aspects of overparameterization (Belkin et al., 2019; Neyshabur et al., 2019) and implicit regularization in gradient descent (Soudry et al., 2018), we still lack a unified understanding of how the structure of parameter space shapes the loss landscape and affects optimization.
@@
This survey proposes parameter space symmetry as a valuable, underexplored perspective for understanding neural networks. Parameter space symmetries are transformations of neural network parameters that leave the output unchanged, e.g. permuting neurons in a hidden layer (Figure 1). One of the most important consequences of these symmetries is that they induces nontrivial structure in the level sets of the loss function. In overparameterized networks, where many parameter configurations can yield the same function output, the loss landscape often contains high dimensional manifolds of global minima (Cooper, 2018). Symmetry accounts for much of this degeneracy by mapping parameters within a level set without changing the network’s function. Understanding these equivalence classes is essential for analyzing generalization, optimization dynamics, and the loss landscape in deep learning.
While symmetry in the data space has been central to geometric deep learning and equivariant models (Bronstein et al., 2021), symmetry in the parameter space has only recently begun to receive sustained attention. Recent work explores parameter symmetry in diverse contexts, including loss landscapes and mode connectivity (Garipov et al., 2018; Draxler et al., 2018), conserved quantities and training dynamics (Kunin et al., 2021; Tanaka & Kunin, 2021), symmetry-based optimization and model averaging (Entezari et al., 2022; Ainsworth et al., 2023), and symmetry-aware sampling in Bayesian inference (Wiese et al., 2023). This survey aims to define symmetries in neural network parameter spaces, unify existing theoretical and algorithmic perspectives, and highlight their relevance to learning theory, optimization, and model analysis. We hope that a clearer understanding of parameter space symmetries will lead to more principled approaches in deep learning theory and practice.
@@ 2.1 Loss-Invariant Parameter Transformations and Symmetry
Let $\Theta$ be the space of parameters and $\mathcal{D}$ be the space of data. In supervised learning, $\mathcal{D}$ is decomposed into $\mathcal{D}_{\text{input}}\times\mathcal{D}_{\text{target}}$, where $\mathcal{D}_{\text{input}}$ and $\mathcal{D}_{\text{target}}$ correspond to the space of inputs and target labels, respectively. Let $f\colon\Theta\times\mathcal{D}_{\text{input}}\to\mathcal{D}_{\text{target}}$ be a neural network function, and $c\colon\mathcal{D}_{\text{target}}\times\mathcal{D}_{\text{target}}\to{\mathbb{R}}$ be a function that measures the discrepancy between the neural network’s prediction and the ground truth label. The loss function $L\colon\Theta\times\mathcal{D}\to{\mathbb{R}}$ is the composition of $f$ and $c$. That is, for $(\theta,(x,y))\in\Theta\times\mathcal{D}$, we define the loss $L(\theta,(x,y)):=c\big(f(\theta,x),\,y\big)$. With occasional exceptions (Bassey et al., 2021; Huang et al., 2018), the parameter space of a neural network is typically a real vector space.
@@
The set of all such transformations forms a group under function composition. To see this, we first examine its structure and properties. Viewing the transformations as functions, we are able to define the composition of two transformations. The composition of two loss-invariant transformations results in another loss-invariant transformation. Since function compositions are associative, the composition of these transformations are associative. Additionally, the identity transformation is always loss preserving, therefore included in our set. Finally, since each transformation is bijective, an inverse exists for every element in the set. These properties of the set of symmetry transformations under composition satisfy the definition of a group.
@@
Following Definition 2.1, the set of all loss-invariant and bijective transformations, which we denote by $G_{\Theta,L}$, forms a group. In practice, we typically consider specific subgroups of $G_{\Theta,L}$, since the full group is often difficult to characterize.
@@ 2.2 Functional Neural Network Symmetry
Parameter space symmetry can be defined in various ways, depending on the transformation preserved, i.e. the neural network function or the loss function, and the scope of the data considered, which can range from all possible data, subsets of data, or a particular distribution (Figure 5).
@@ 2.2.1 Examples: Symmetries in Common Neural Network Components
Often appearing in equivariant neural networks (Weiler et al., 2018; Weiler & Cesa, 2019), a radial rescaling activation $\sigma\colon{\mathbb{R}}^{h}\to{\mathbb{R}}^{h}$ has the form $\sigma(z)=f(\|z\|)z$ for some function $f\colon{\mathbb{R}}\to{\mathbb{R}}$. A two-layer neural network $f(W_{2},W_{1},X)=W_{2}\sigma(W_{1}X)$ with a radial rescaling activation $\sigma$ has an $\text{O}_{h}({\mathbb{R}})$ symmetry, acting on $\Theta$ by $g\cdot(W_{2},W_{1})=(W_{2}g^{-1},gW_{1})$, for $g\in\text{O}_{h}({\mathbb{R}})$.
@@
Example 2.6 (Batch Normalization (Ioffe & Szegedy, 2015)).
Batchnorm standardizes inputs of a layer across a mini-batch: $\text{BN}(Z)=\frac{Z-\text{E}[Z]}{\sqrt{\text{Var}[Z]}}$, where the expectation and variance are computed row-wise over $Z$. Assuming $Z$ is a linear feature $Z=WX$ where $W\in\Theta={\mathbb{R}}^{m\times n}$ and $X\in{\mathbb{R}}^{n\times k}$, then the batchnorm function has a positive scaling symmetry. Batchnorm is equivariant with respect to the group ${\mathbb{R}}_{>0}^{m}$, which consists of diagonal matrix with positive diagonal entries, acting on $\Theta$ by $g\cdot W=gW$, for $g\in{\mathbb{R}}_{>0}^{m}$ (Kunin et al., 2021). Other normalization layers, such as layer normalization (Ba et al., 2016), group normalization (Wu & He, 2018), and weight normalization (Salimans & Kingma, 2016), often have scaling symmetry of similar forms.
@@
Many architectures have parameter symmetries given by a symmetric group. This symmetry can often be interpreted as permutation of neurons or components along with their associated weights, as shown in Examples 2.7, 2.8 and Figure 3(c). Similar forms of symmetry have also been identified in recurrent neural networks (Albertini et al., 1995).
@@
Example 2.9 (Radial basis function networks (Broomhead & Lowe, 1988)).
@@ 2.2.2 Example: Symmetries in Transformers
2.3 Relaxed Definitions of Symmetry
@@ 2.4 Parameter Identifiability and Completeness of Symmetry
Having introduced a range of examples highlighting that symmetry is ubiquitous in neural network parameter spaces, we now turn to the fundamental question of whether known parameter symmetries account for all the ways in which parameters can represent the same function. This question concerns parameter identifiability—the extent to which a neural network’s function determines its parameters, which is important in interpreting and comparing trained networks (Ran & Hu, 2017). This section formalizes the relationship between symmetry and identifiability using the realization map from parameters to functions, and surveys results on when known symmetries are sufficient to characterize all such redundancy.
Formally, we consider the realization map $\rho\colon\Theta\to\mathcal{F},\text{where }\mathcal{F}=\{f\colon\mathcal{D}_{\text{input}}\to\mathcal{D}_{\text{output}}\}$, which associates each parameter $\theta\in\Theta$ with the function $\rho(\theta)\in\mathcal{F}$ it defines (Grigsby et al., 2025). The fiber of a function $f$ under $\rho$, denoted $\rho^{-1}(f):=\{\theta\in\Theta~|~\rho(\theta)=f\}$, is the set of all parameters that map to the same function $f$. In many architectures, $\rho$ is not injective, and fibers contain multiple distinct parameters that define the same function. This implies that parameters are, in general, not uniquely determined by the function alone.
@@
For several architectures, such completeness has been established. For example, Gaussian radial basis function networks are known to be identifiable up to permutation symmetry, indicating that permutation symmetry is complete for these architectures (Kůrková & Neruda, 1994). For feedforward networks with tanh activations, if two networks represent the same function, then they have the same architecture and parameters up to permutation and sign-flip symmetries, under a known set of conditions (Sussmann, 1992; Fefferman & Markel, 1993). Similar completeness results have also been established for complex-valued tanh networks (Nitta, 2003; Kobayashi, 2010), tanh recurrent neural networks (Kimura, 2002), networks with activation functions $\sigma$ satisfying $\sigma(0)=0,\sigma^{\prime}(0)\neq 0$, and $\sigma^{\prime\prime}(0)=0$ (Albertini & Sontag, 1993), and networks with asymptotically constant activation functions (Kůrková & Kainen, 1994). For Mixture-of-Experts (MoE) architectures, permutation-translation symmetries of the gating parameters were recently proven to fully characterize functional equivalence (Tran et al., 2025b).
@@
The identifiability of ReLU networks is less straightforward. In particular, ReLU networks can contain hidden symmetries—transformations that preserve the realized function but not captured by permutation and positive scaling symmetries. Grigsby et al. (2023) describe several mechanisms through which hidden symmetries can arise. For example, for a single ReLU neuron with scalar input $x$ and realization map $\rho\colon(a,b)\mapsto(x\mapsto\text{ReLU}(ax+b))$, the fiber of the constant function 0, $\rho^{-1}(0)$, contains $(0,0)$ and (0,1)(0,-1), which are not related by rescaling and permutation (Grigsby et al., 2025). To formalize this variability, Grigsby et al. (2025) introduce the functional dimension, defined as the rank of the Jacobian of the realization map $\rho$, which reflects the local dimensionality of the set of functions realized near a given parameter. They show that the functional dimension may not be constant even within the same fiber. This suggest a complicated structure of symmetries in ReLU networks.
@@
Other recent work investigates when ReLU networks are identifiable up to positive scaling and permutation symmetries. (Petzka et al., 2020) prove that excluding a degenerate case where two neurons have identical zero hyperplanes, two-layer ReLU networks do not have additional symmetries besides permutation and positive scaling. (Rolnick & Kording, 2020) prove that under the assumption that networks within each linear region compute the same linear function, boundaries between linear regions uniquely determine parameters up to permutation and positive scaling. For ReLU networks with either non-increasing widths (Bui Thi Mai & Lampert, 2020) or all layers wider than input (Grigsby et al., 2023), a positive measure subset of the parameters have no symmetries besides permutation and positive scaling (hidden symmetries). Grigsby et al. (2023) also empirically show that the probability of having no hidden symmetry increases with input and layer width and decreases with depth. (Bona-Pellissier et al., 2023) provide sufficient conditions under which two ReLU network are identical up to permutation and positive scaling if their output agrees on a subset of the input space.
@@
The identifiability of neural networks that are polynomial in their parameters can often be analyzed using tools from algebraic geometry (Marchetti et al., 2025). A notable example is multi-layer perceptrons with polynomial activation functions, or polynomial neural networks. Kileel et al. (2019) conjecture that permutation and scaling are the only symmetries in generic polynomial neural networks, which was later proved for the monomial case (Finkel et al., 2025). For more general polynomial activation functions, there are only finitely many data-independent parameter symmetries Shahverdi et al. (2025). Convolutional neural networks with polynomial activations generically have no nontrivial data-independent parameter symmetries. Another example of a polynomial function is lightening self-attention $(W^{Q},W^{K},W^{V})\mapsto QW^{Q}(KW^{K})^{T}VW^{V}$, where the softmax normalization is omitted from the standard attention mechanism (Example 2.10). Henry et al. (2025) prove that in networks composed of lightning self- attention layers, the only symmetries are (1) scaling $W_{Q}W_{K}^{T}$ and $W_{V}$ by a constant, (2) scaling $W_{Q}$ and $W_{K}$ by an invertible matrix, and (3) scaling $W_{V}$ of one layer and $W_{Q},W_{K}$ in the next layer by an invertible matrix. They conjecture and verify numerically that adding back the softmax breaks only the first of these symmetries.
@@ 3.3 Removing Symmetry: Model Compression and Reduced Search Space
While symmetry enriches the loss landscape by introducing structure and multiplicity, removing symmetry reveals a simpler geometry—one that is more compact to represent, and often more amenable to analysis. Symmetry removal effectively collapses each symmetry orbit to a canonical representative (Sorensen, 2020). This operation alters the shape of the loss landscape—not by changing its values, but by restricting attention to a lower-dimensional quotient space where each point represents a unique function. As a result, the loss landscape becomes less redundant, more compact, and easier to sample from and optimize.
One major motivation for removing symmetry is to reduce the dimensionality of the parameter space without altering the function space. This operation, often known as model compression, is of critical importance in lowering storage cost, improving inference speed, and efficient deployment of machine learning models. Every symmetry implies a set of directions in which the loss is constant, stretching level sets into high-dimensional manifolds. Quotienting out these directions yields a more compact version of the loss landscape, where each point corresponds to a unique function. In the context of model compression, this means we can represent networks more efficiently. For example, Ganev et al. (2022) compress radial neural networks by factoring out orthogonal symmetries, and Sourek et al. (2021) exploit structural symmetries in computational graphs to merge redundant nodes. These methods reduce storage costs and inference time, without affecting expressivity or accuracy.
@@
In Bayesian neural networks, removing symmetry reduces the effective search space by collapsing functionally equivalent modes in the posterior (Lim et al., 2024b). In high-dimensional models like neural networks, symmetries can cause posterior distributions to become multi-modal in a functionally redundant way. This impairs MCMC mixing and complicates interpretation. By mapping parameters to a canonical representative—effectively quotienting out permutation or sign-flip symmetries—Wiese et al. (2023) and Laurent et al. (2024) show that it is possible to explore a more informative and diverse set of functionally distinct solutions. (Xiao et al., 2023) show that removing permutation symmetry leads to a compact representation that helps directly compare trained Bayesian neural networks across sampling methods.
Finally, removing symmetry alters the curvature and critical point structure of the landscape, with direct implications for optimization. Symmetry-induced flat directions often give rise to plateaus or saddle points that slow down training. By projecting the loss landscape onto a symmetry-reduced space, these degenerate directions are removed, yielding a more strongly convex surface. Leake & Vishnoi (2021) show that continuous symmetries can be used to construct convex polytopes, allowing nonconvex optimization problems to be reformulated as linear optimization problems over polytopes, which are computationally more tractable.